# Intensity-Efficient and Intensity-Positional Allocations

### 01 Mar 2026 v6

@@ 2 Measurement-Theoretic Foundations
In line with these interpretations and the literature, agent $l$’s binary preference relation $\succsim_{l}$ on $A$ is derived from their intensity relation $\dot{\succsim}_{l}$ by $a\succsim_{l}b\Leftrightarrow(a,b)\,\dot{\succsim}_{l}\,(b,a)$. We will refer to $\succsim_{l}$ as the preference relation that is induced by $\dot{\succsim}_{l}$. As usual, the asymmetric and symmetric parts of $\succsim_{l}$ are denoted by $\succ_{l}$ and $\sim_{l}$. Under the structure that $\dot{\succsim}_{l}$ will be endowed with below, both this relation and its induced $\succsim_{l}$ are weak orders on their respective domains.
@@
The definitions that follow omit universal quantifiers, yet they should be understood as imposing conditions that apply to all objects in the respective domains.
Definition 1.
@@
Scott (1964), completing and extending the analysis that was initiated in Scott and Suppes (1958), characterized the existence of a utility-difference representation over a finite set by means of the following three axioms on $\dot{\succsim}_{l}$:
$(\mathbf{1_{\mathit{D}}})$ [Completeness]. $(a,b)\,\dot{\succsim}_{l}\,(c,d)$ or $(c,d)\,\dot{\succsim}_{l}\,(a,b)$.
$(\mathbf{2_{\mathit{D}}})$ [Cancellation]. $(b_{i},c_{i})\,\dot{\succsim}_{l}\,(b_{\pi(i)},c_{\sigma(i)})$ for all $i<t$, $i>0$, implies $(b_{\pi(0)},c_{\sigma(0)})\,\dot{\succsim}_{l}\,(b_{0},c_{0})$ for all sequences $b_{0},b_{1},\ldots,b_{t-1},c_{0},\ldots,c_{t-1}\in A$ and all permutations $\pi,\sigma$ of $\{0,\ldots,t-1\}$, where $t>0$.
$(\mathbf{3_{\mathit{D}}})$ [Reversal]. $(a,b)\,\dot{\succsim}_{l}\,(c,d)$ implies $(d,c)\,\dot{\succsim}_{l}\,(b,a)$.
The first imposes universal comparability on $\dot{\succsim}_{l}$. The third is essential for $\dot{\succsim}_{l}$ to be interpreted as a preference intensity relation rather than as an arbitrary preference relation over pairs. The second is known to have “no simple way of summarizing what it says about choice behavior” (Luce and Suppes, 1965, p. 277). Scott (1964, p. 244-5) explained how $(\mathbf{2_{\mathit{D}}})$ implies transitivity of $\dot{\succsim}_{l}$, but also remarked that $(\mathbf{2_{\mathit{D}}})$ is an infinite bundle of conditions (for each $t>0$, each $\pi,\sigma$), and it was shown in Scott and Suppes (1958) that no finite number of them could be sufficient”. Once the number $n$ of alternatives in $A$ is fixed, there are obviously finitely many such conditions to be satisfied. Scott’s remark, however, clarifies that it is impossible to reduce this condition to a finite number of statements/axioms that, together with $(\mathbf{1_{\mathit{D}}})$ and $(\mathbf{3_{\mathit{D}}})$, would be sufficient to characterize (3) for any finite $n$.
In Gerasimou (2021, 2022) this author recently proposed and characterized the class of ordinal intensity relations that admit the following more general numerical representation.
Definition 2.
@@
with a strict inequality on the left part of (6) implying a strict inequality on the right.
In words, $\widehat{s}_{l}$ represents the agent’s intensity weak ordering [(4)] and the preferences induced by it [(5)], further ensuring that these preferences are themselves weakly ordered and intensities are monotonically increasing in the preference ordering [(6)]. Such a representation is possible if and only if $\dot{\succsim}_{l}$ is a weak order on $A\times A$ that additionally satisfies $(\mathbf{3_{\mathit{D}}})$ and the following condition, which states that $a$ is preferred to $b$ if and only if the intensity difference between $a$ and $c$ is higher than that between $b$ and $c$.
$(\mathbf{4_{\mathit{D}}})$ [Translocation Consistency]. $(a,c)\,\dot{\succsim}_{l}\,(b,c)$ if and only if $(a,b)\,\dot{\succsim}_{l}\,(b,a)$.
We denote by $\mathcal{I}$ the class of intensity relations that admit this ordinal representation.
@@
A relation $\dot{\succsim}_{l}$ that is represented by $\widehat{s}_{l}$ is utility-difference representable if and only if some strictly increasing transformation $\widetilde{s}_{l}$ of $\widehat{s}_{l}$ satisfies
@@
The proof of this and most other results appear in the Appendix.
The normalization of $\widehat{s}_{l}$ introduced next is important for the rest of this paper.
Definition 3.
A preference intensity function $\widehat{s}_{l}$ represents $\dot{\succsim}_{l}$ canonically if
@@
with $k$ the number of intensity equivalence classes $a_{i},a_{j}]\in(A\times A)_{\dot{\sim}_{l}}$ such that $a_{i}\succ_{l}a_{j}$.
We will denote by $s_{l}$ the canonical intensity function that represents $\dot{\succsim}_{l}$. The usefulness of this normalization for our purposes lies in the fact that the integer it associates with a pair reveals the rank position of that pair in the agent’s intensity ordering.
@@
$(\mathbf{5_{\mathit{D}}})$ [Strictness]. $(a,b)\,\mathord{\not\mathrel{\dot{\sim}_{l}}}\,(c,d)$ for all distinct off-diagonal pairs $(a,b),(c,d)\in A\times A$.
This condition rules out the possibility of the agent preferring $a$ to $b$ exactly as much as they prefer $c$ to $d$ for distinct pairs of distinct alternatives $(a,b)$ and $(c,d)$. It is therefore analogous to the standard preference-strictness postulate which, in fact, it implies. As we elaborate in Section 3, strict intensities are useful for our purposes because they ensure that the canonical intensity functions of all agents are onto the same set $\{-k,\ldots,-2,-1,0,1,2,\ldots,k\}$, where $k\equiv{n\choose 2}$. In particular, the only non-trivial intensity-equivalence class when $\dot{\succsim}_{l}$ satisfies $(\mathbf{5_{\mathit{D}}})$ is the diagonal subset of $A\times A$, with every pair $(a,a)$ in this set being mapped to 0 by the canonical representation of $\dot{\succsim}_{l}$. For illustration, the novel computational output in columns 3–4 of Table 1 lists the number of distinct strict intensity orderings that admit,respectively, a (pseudo-cardinal) utility-difference representation and an ordinal representation when $n\leq 5$.
Table 1: The number of distinct intensity relations of different classes for small values of $n$.
Strict & utility-difference Strict Strict ordinal
Strict representable ordinally representable intensity
$n$ orders intensity relations intensity relations profiles
3 6 12 12 $12^{3}$
4 24 240 384 $384^{4}$
5 120 13,680 92,160 $92,160^{5}$
Source: author’s computations with the Minizinc constraint-programming tool (Nethercote et al., 2007).
We conclude this section by commenting on possible sources of intensity information. Section 2 in Gerasimou (2021) discusses various non-choice variables that, according to different literatures, might be thought of as having intensity-revealing potential. Prominent examples include willingness to pay; 50/50 lottery choices; survey-based ratings; and response times. Section 5 of that paper, moreover, shows how such data can be analysed alongside choices by means of otherwise standard revealed-preference theoretic techniques, providing testable foundations for the hypothesis that an agent’s overall behaviour complies with utility maximization and ordinal intensity revelation. The present paper, in contrast, operates within the environment of the allocation problem, where agents are traditionally assumed to submit their full ordinal or cardinal preference reports directly. This is different from a lab, field or other environment that might, in principle, provide the analyst with agents’ choices at several menus at different points in time, together with intensity-revealing information at those menus. As we elaborate formally in Section 6, for this specific environment we assume that agents’ ordinal preferences are available, and that their ordinal intensities are elicited by asking them directly—and conditional on their submitted preferences—the kinds of simple binary questions that were mentioned earlier.
@@ 1 $(a_{1},a_{2})$ $(a_{6},a_{1})$ $(a_{6},a_{5})$ $(a_{3},a_{4})$ $(a_{4},a_{1})$ $(a_{1},a_{3})$
For $n=4$ let the envy-free and top-preference-coincident sets that partition $N$ and $A$ be defined by $(N^{1},A^{1})=(\{v_{4}\},\{a_{4}\})$ and $(N^{2},A^{2})=(\{v_{1},v_{2},v_{3}\},\{a_{1},a_{2},a_{3}\})$, as in Table 4. The Pareto-efficient allocations are $p:=(a_{1},a_{2},a_{3},a_{4})$, $q:=(a_{1},a_{3},a_{2},a_{4})$, $r:=(a_{2},a_{3},a_{1},a_{4})$, $x:=(a_{2},a_{1},a_{3},a_{4})$, $y:=(a_{3},a_{1},a_{2},a_{4})$ and $z:=(a_{3},a_{2},a_{1},a_{4})$. Among them, $r$ is intensity-efficient: it dominates $x$; is incomparable to $p$ and $y$; and is equivalent to $q$, $z$ (these are dominated by $p$, $y$, respectively). Yet $N^{2}$ is (non-strictly) monotonic but not homeo-monotonic with respect to $A^{2}$: $s_{1}(a_{1},a_{2})=s_{1}(a_{1},a_{2})>s_{3}(a_{1},a_{2})$ and $s_{1}(a_{2},a_{3})=s_{2}(a_{2},a_{3})>s_{3}(a_{2},a_{3})$, but $s_{1}(a_{1},a_{3})=s_{2}(a_{1},a_{3})<s_{3}(a_{1},a_{3})$. $\blacklozenge$
@@ 5.2 Intensity-Positional Allocations
Analogously, a Borda allocation at profile $S\in\mathcal{\widehat{S}}$—henceforth denoted by $\mathsf{Bo}(S)\equiv\mathsf{Bo}(P_{S})$—is defined as in (14) but with the Borda scores $b_{l}^{P}$ at the strict preference profile $P_{S}$ replacing the intensity-rank scores $r_{l}^{S}$ at the strict intensity profile $S$.
@@
An intensity-positional allocation is a Borda allocation (hence Pareto-efficient): for all $S\in\mathcal{\widehat{S}}$, $\mathsf{IP}(S)\subseteq\mathsf{Bo}(S)$. The converse is false.
In other words, Borda allocations—which, as is well-known, are Pareto-efficient—are refined by the intensity-positional criterion. That this refinement is generally strict can in turn be verified with the following $n=3$ example profile:
$(a,c)\dot{\succ}_{1}(a,b)\dot{\succ}_{1}(b,c)$; $(a,c)\dot{\succ}_{2}(b,c)\dot{\succ}_{2}(a,b)$; $(c,a)\dot{\succ}_{3}(c,b)\dot{\succ}_{3}(b,a)$.
The reader can readily check that both Pareto-efficient allocations $(a,b,c)$ and $(b,a,c)$ are selected at this profile under $\mathsf{Bo}$, whereas only the former is so under $\mathsf{IP}$ (and that this allocation is also intensity-efficient). This is intuitive: the first two agents have identical preferences, but the intensity rank of $b$ is higher for the second agent, reflecting the fact that preference intensity between $a$ and $b$ is higher for agent 1 than it is for agent 2. The intensity-positional criterion accounts for this interpersonal difference, whereas the Pareto and Borda criteria do not.
In this example it so happened that the set of intensity-positional and intensity-efficient allocations coincided. Yet these criteria are distinct in general.
@@
The sets of intensity-efficient and intensity-positional allocations at a profile $S\in\mathcal{\widehat{S}}$ are generally distinct.
The first implication of this claim—namely, that an intensity-positional allocation may not be intensity-efficient—has already been established in Example 5. The converse implication can also be proved by modifying the profile of Example 3 (Table 4) so that $s_{2}(a_{1},a_{3})=5$, $s_{2}(a_{2},a_{4})=4$ and $s_{3}(a_{1},a_{2})=4$, $s_{3}(a_{2},a_{4})=3$, $s_{3}(a_{2},a_{3})=2$, $s_{3}(a_{3},a_{4})=1$. In this case, the set of intensity-efficient allocations contains three elements, one of which—namely, $(a_{3},a_{2},a_{1},a_{4})$—is not intensity-positional.
@@
6 Eliciting Ordinal Intensities and Improving Upon Pareto-Efficient Allocations
We conclude with an illustration of the potential usefulness of the preceding analysis in matching-theoretic problems. In particular, we study a matching rule/mechanism that takes agents’ ordinal intensities alongside their ordinal preferences as its input and produces an allocation that improves upon Pareto-efficiency, either in the intensity-dominating or intensity-positional direction.
Before doing so, we clarify how ordinal intensities can be elicited alongside preferences in the first place in order for such a mechanism to become operational. To this end, we show below how the properties of intensity relations that are representable by ordinal preference intensity functions can simplify this elicitation problem very considerably.
Proposition 5.
Given a strict preference order $\succ_{l}$, an intensity relation $\dot{\succsim}_{l}\in\mathcal$$\widehat${I}} that induces $\succ_{l}$ is elicited in at most (n1)(n+1)(n-1)(n+1) steps, hence in $O(n^{2})$ time.
Turning now to the problem of using as input a strict intensity profile SS—and the strict preference profile $P_{S}$ nested within it—to find an allocation that improves upon Pareto-efficiency, we proceed by recalling first some important landmarks from the matching-theoretic literature with one-sided preferences where no intensity information is available. The study of such rules goes back to Shapley and Scarf (1974) and David Gale’s Top-Trading Cycles (TTC) algorithm. The premise there is that each agent has an initial endowment from the set of items that will be (re-)assigned once the algorithm is run. TTC then operates sequentially by implementing swaps between these initial assignments in accordance with the agents’ preferences, giving priority to an item’s current assignee in situations where additional agents prefer this over their own endowment. Shapley and Scarf (1974) proved that TTC leads to a core allocation. Roth and Postlewaite (1977) later showed that this core allocation, identifiable by TTC, is unique. Extending the results in Roth (1982), Ma (1994) established that TTC is the unique rule that leads to a Pareto-efficient allocation and, in addition, satisfies strategy-proofness and individual-rationality.
Hylland and Zeckhauser (1979) introduced randomness to study the assignment problem without initial endowments and, assuming that all agents have an equal amount of some artificial currency and expected-utility preferences over lotteries over allocations, studied that which is associated with (or approximated by) Competitive Equilibrium from Equal Incomes. This mechanism is not strategy-proof. Also introducing randomness, but without assuming expected-utility preferences or employing the competitive-equilibrium apparatus, Abdulkadiroğlu and Sönmez (1998) showed that TTC with random initial endowments is equivalent to the Random Priority (RP) mechanism. The latter does not involve initial endowments; ranks agents under a random priority order; and lets them choose, accordingly, their most preferred item from those still available. These authors also clarified that a deterministic matching mechanism leads to a Pareto-efficient allocation if and only if is a Random Priority mechanism.
Despite the ex post Pareto-efficiency of RP allocations, hence of core/TTC allocations under random endowments, Zhou (1990) showed that these are generally ex ante inefficient. Bogomolnaia and Moulin (2001), finally, demonstrated that they are also ordinally inefficient in the sense that they may be dominated by another random allocation under the stochastic-dominance relation that is defined by all agents’ preferences. In response to this realization, the latter authors introduced ordinally-efficient allocations as those that are not first-order stochastically dominated by any other random allocation, and proved that their Probabilistic Serial (PS) mechanism, although not strategy-proof, leads to an ordinally-efficient allocation with desirable—under expected-utility preferences over allocations—ex ante fairness properties.
Informed by the main insights of the above-cited literature, we now consider the following algorithm for the problem at hand:
Input: A strict intensity profile $S\in\mathcal{\widehat{S}}$.
Step 1: Choose a linear order $\gg$ over $N$ from the uniform distribution with support $n!$.
Step 2: Apply RP under $\gg$ on $P_{S}$ to arrive at a Pareto-efficient allocation $x$.
Step 3: Reorder the entries in $x$ according to $\gg$, with $j>i$ for $x_{i},x_{j}$ if and only if $i\gg j$.
Step 4: Search for the $\gg$-lowest agent $v_{j}\in N\setminus\{v_{1}\}$ such that $s_{1}(x_{j},x_{1})>s_{j}(x_{j},x_{1})$. If such an agent exists, swap $x_{1}$ with $x_{j}$ and remove $v_{1},v_{j}$ from $N$ and $x_{1},x_{j}$ from $A$.
If no such agent exists, remove $v_{1}$, $x_{1}$.
Repeat Step 4 until no agents remain.
The first two steps setup the RP algorithm. The third one merely relabels the items in the allocation that was arrived at earlier, so that the first (second, third, …) item belongs to the first-(second-, third-, …)ranked agent in the ordering that was drawn at the beginning. The fourth step is the essential new addition. It starts with agent 1 and searches—without loss, from the bottom up—for an agent $j$ whose RP assignment, $x_{j}$, is preferred by agent 1 to their own assignment, $x_{1}$, more than it is so by agent $j$. If such an agent exists, the algorithm implements the swap, removes the two agents and the respective items, then repeats this step with the remaining agents and items. If not, then agent 1 and $x_{1}$ are removed before the step is re-applied. This forced exit is important in two ways. First, it prevents potentially indefinite loops, for example in cases where the RP allocation is already intensity-efficient. Second, as is clarified below, it ensures that the computational complexity of the whole process falls in the manageable quadratic class.
Proposition 6.
If $S\in\mathcal{\widehat{S}}$, and either $n=3$ or $S$ is effectively homeo-monotonic, then the algorithm ends in at most ${n}\choose{2}$} steps—hence in $O(n^{2})$ time—and either produces the Pareto-efficient allocation of its RP-implementing step or one that intensity-dominates it.
The allocation delivered by this algorithm is not intensity-efficient in general. This is so because the forced exit of agents and items in the fourth step means that not all pairwise possibilities for intensity-dominance-driven improvements are considered in the transition from the RP-mandated Pareto-efficient allocation in Step 2 to the final allocation. Clearly, however, the algorithm improves upon RP and Pareto-efficiency—even if partially so—in the direction of intensity-efficiency, without adding intractable computational complexity relative to RP. It is presently unknown if this algorithm can be modified to deliver an intensity-efficient allocation—whenever one is known to exist—without requiring factorial search over all possible priority orderings over agents in Step 2.555Suppose, instead, that an intensity-positional allocation is the welfare-improving criterion of interest. Such an allocation exists for any strict intensity profile and can be identified via simple summation once agents’ ordinal intensities are elicited. Owing to these facts, it is a straightforward implication of the discussion around Proposition 6 that, in order for the algorithm’s RP allocation to be converted into an intensity-positional one, it suffices to replace the last two steps by any deterministic or random criterion that selects one among the possibly many intensity-positional allocations corresponding to the given profile. Moreover, by Proposition 5 and the linear complexity of the RP algorithm (Manlove, 2013, Section 6.2), this process terminates in at most (n1)(n+1)(n-1)(n+1) as well.
In light of the discussion preceding the statement of Proposition 6, it is clear that RP can be replaced by PS in Step 2, depending on the desired resolution between strategy-proofness and ordinal-efficiency. Regardless of whether RP, PS or some other mechanism is used to elicit the preference component of the input intensity profile in the manner dictated in the proof of Proposition 5, however, one is naturally interested to know if, when operating under the matching rule defined by the above algorithm, agents are incentivized to report their intensities truthfully. The answer is negative. To see this, suppose $n=3$ and let $S\in\mathcal{\widehat{S}}$ be such that $a\succ_{1}b\succ_{1}c$; $a\succ_{2}b\succ_{2}c$; $c\succ_{3}b\succ_{3}a$; $(a,b)\dot{\succ}_{1}(b,c)$; $(b,c)\dot{\succ}_{2}(a,b)$; $(c,b)\dot{\succ}_{3}(b,a)$. The possible orderings over $N$ are: 1. $v_{1}\gg v_{2}\gg v_{3}$; 2. $v_{1}\gg v_{3}\gg v_{2}$; 3. $v_{2}\gg v_{1}\gg v_{3}$; 4. $v_{2}\gg v_{3}\gg v_{1}$; 5. $v_{3}\gg v_{1}\gg v_{2}$; 6. $v_{3}\gg v_{2}\gg v_{1}$. The two Pareto-efficient allocations are $(a,b,c)$, $(b,a,c)$, and the former intensity-dominates the latter. Suppose the algorithm’s second step delivers this intensity-efficient allocation (under RP, this happens with prob. $1\over 2$}). If intensities are reported truthfully, there is no order $\gg$ under which the algorithm’s subsequent steps alter it. But if $(b,c)\dot{\succ}^{\prime}_{2}(a,b)$ is the second agent’s misreported intensity, then under the 3rd, 4th and 6th $\gg$-orders (i.e. with prob. $1\over 2$}) the algorithm reassigns $b$ to $v_{1}$ and $a$ to $v_{2}$, making the latter agent better off.
Interestingly, the incentives to misreport one’s intensities in this way, namely toward increasing one’s chances of getting their most preferred alternative, act in the opposite direction to the incentives toward misreporting one’s preferences in the Boston/Immediate-Acceptance mechanism of the distinct two-sided (School Choice) matching problem. Specifically, as has been discussed by Pathak and Sönmez (2008); Abdulkadiroğlu, Che, and Yasuda (2011) and other authors, under this mechanism students may (and often do) submit their second- or third-ranked school as their supposedly most preferred one if their actual top-ranked school is very popular, out of concern that in the event they are not assigned to it they may end up at one of their lower-ranked schools if their second or third most preferred ones have in the meantime filled up. Abdulkadiroğlu et al. (2011), however, followed by Pycia (2011), Featherstone and Niederle (2016) and other studies, argued that this drawback on incentive grounds of the Boston mechanism may be outweighed by the total welfare gains associated with such misreporting once the students’ intensities are accounted for. In line with the prior relevant literature, intensity comparisons in that paper’s Bayesian-Nash equilibrium analysis are derived from intra- and inter-personal differences in cardinal utility functions that emerge from expected-utility preferences over lotteries over schools. In that framework too, finally, the intensity information that is conveyed to the social planner is assumed to exist without having been elicited incentive-compatibly, and it is unclear if such elicitation is possible. In relation to this, Pycia and Ünver (2025) recently showed that, in environments without monetary transfers such as the one considered above, only ordinal mechanisms are group strategy-proof.

### 14 Apr 2026 v8

@@ 2 Preliminaries
In line with these interpretations, and with the literature, agent $l$’s binary preference relation $\succsim_{l}$ on $A$ is derived from their intensity relation $\dot{\succsim}_{l}$ by $a\succsim_{l}b\Leftrightarrow(a,b)\,\dot{\succsim}_{l}\,(b,a)$. We will refer to $\succsim_{l}$ as the preference relation that is induced by $\dot{\succsim}_{l}$. As usual, the asymmetric and symmetric parts of $\succsim_{l}$ are denoted by $\succ_{l}$ and $\sim_{l}$. Under the structure that $\dot{\succsim}_{l}$ will be endowed with below, both this relation and its induced $\succsim_{l}$ are weak orders on their respective domains.
@@
with a strict inequality on the left part of (6) implying a strict inequality on the right. In words, $\widehat{s}_{l}$ represents the agent’s intensity weak ordering [(4)] and the preferences induced by it [(5)], further ensuring that these preferences are themselves weakly ordered and intensities are monotonically increasing in the preference ordering [(6)]. We denote by $\mathcal{I}$ the class of intensity relations that admit this ordinal representation.
@@
A relation $\dot{\succsim}_{l}$ that is represented by a preference intensity function $\widehat{s}_{l}$ is utility-difference representable if and only if a strictly increasing transformation $\widetilde{s}_{l}$ of $\widehat{s}_{l}$ satisfies
@@
The proof of this and most other results appears in the Appendix.
Next, we will say that a preference intensity function $\widehat{s}_{l}$ represents $\dot{\succsim}_{l}$ canonically if
@@
where $k$ is the number of intensity equivalence classes $a_{i},a_{j}]\in(A\times A)_{\dot{\sim}_{l}}$ such that $a_{i}\succ_{l}a_{j}$. We will denote by $s_{l}$ the canonical intensity function that represents $\dot{\succsim}_{l}$. As will be clarified later, the usefulness of this normalization for our purposes lies in the fact that the integer value it associates with a pair reveals the rank position of that pair in the agent’s intensity ordering.
@@
Strictness. $(a,b)\,\mathord{\not\mathrel{\dot{\sim}_{l}}}\,(c,d)$ for all distinct off-diagonal pairs of items $(a,b)$ and $(c,d)$.
This condition rules out the possibility of the agent preferring $a$ to $b$ exactly as much as they prefer $c$ to $d$ for distinct pairs of distinct alternatives $(a,b)$ and $(c,d)$. It is therefore analogous to the standard preference-strictness postulate which, in fact, it implies. As we elaborate in Section 3, strict intensities are useful for our purposes because they ensure that the canonical intensity functions of all agents are onto the same set $\{-k,\ldots,-2,-1,0,1,2,\ldots,k\}$, where $k\equiv{n\choose 2}$. In particular, the only non-trivial intensity-equivalence class when $\dot{\succsim}_{l}$ satisfies Strictness is the diagonal subset of $A\times A$, with every pair $(a,a)$ in this set being mapped to 0 by the canonical representation of $\dot{\succsim}_{l}$.
We conclude this section by commenting on possible sources of intensity information. Section 2 in Gerasimou (2021) discusses various non-choice variables that, according to different literatures, might be thought of as having intensity-revealing potential. Prominent examples include willingness to pay; 50/50 lottery choices; survey-based ratings; and response times. Section 5 of that paper, moreover, shows how such data can be analysed alongside choices by means of otherwise standard revealed-preference theoretic techniques, providing testable foundations for the hypothesis that an agent’s overall behaviour complies with utility maximization and ordinal intensity revelation. The present paper, in contrast, operates within the environment of the allocation problem, where agents are traditionally assumed to submit their full ordinal or cardinal preference reports directly. This is different from a lab, field or other environment that might, in principle, provide the analyst with agents’ choices at several menus at different points in time, together with intensity-revealing information at those menus. For this specific environment we assume that agents’ ordinal preferences are available, and that their ordinal intensities are elicited by asking them directly—and conditional on their submitted preferences—the kinds of simple binary questions that were mentioned earlier and which pertain to ranking preference improvements.
@@ 1 $(a_{1},a_{2})$ $(a_{6},a_{1})$ $(a_{6},a_{5})$ $(a_{3},a_{4})$ $(a_{4},a_{1})$ $(a_{1},a_{3})$
For $n=4$ let the envy-free and top-preference-coincident sets that partition $N$ and $A$ be defined by $(N^{1},A^{1})=(\{v_{4}\},\{a_{4}\})$ and $(N^{2},A^{2})=(\{v_{1},v_{2},v_{3}\},\{a_{1},a_{2},a_{3}\})$, as in Table 3. The Pareto-efficient allocations are $p:=(a_{1},a_{2},a_{3},a_{4})$, $q:=(a_{1},a_{3},a_{2},a_{4})$, $r:=(a_{2},a_{3},a_{1},a_{4})$, $x:=(a_{2},a_{1},a_{3},a_{4})$, $y:=(a_{3},a_{1},a_{2},a_{4})$ and $z:=(a_{3},a_{2},a_{1},a_{4})$. Among them, $r$ is intensity-efficient: it dominates $x$; is incomparable to $p$ and $y$; and is equivalent to $q$, $z$ (these are dominated by $p$, $y$, respectively). Yet $N^{2}$ is (non-strictly) monotonic but not homeo-monotonic with respect to $A^{2}$: $s_{1}(a_{1},a_{2})=s_{2}(a_{1},a_{2})>s_{3}(a_{1},a_{2})$ and $s_{1}(a_{2},a_{3})=s_{2}(a_{2},a_{3})>s_{3}(a_{2},a_{3})$, but $s_{1}(a_{1},a_{3})=s_{2}(a_{1},a_{3})<s_{3}(a_{1},a_{3})$. $\blacklozenge$
@@ 5.2 Intensity-Positional Allocations
An intensity-positional allocation maximizes the total of agents’ ordinal intensity-rank scores at their assigned items, and thus resolves trade-offs across agents globally rather than only through pairwise dominance comparisons. Indeed, as was noted earlier, an item’s intensity-rank score contains information about the number of alternatives that are ranked below it as well as about the relative intensity difference between that item and the one ranked at the bottom. This double channel—and trade-offs thereof—guides the intensity-positional allocation criterion and makes it distinct from Borda’s criterion that is guided by the former channel only.
By analogy, a Borda allocation at profile $S\in\mathcal{\widehat{S}}$—henceforth denoted by $\mathsf{Bo}(S)\equiv\mathsf{Bo}(P_{S})$—is defined as in (11) but with the Borda scores $b_{l}^{P}$ at the strict preference profile $P_{S}$ replacing the intensity-rank scores $r_{l}^{S}$ at the strict intensity profile $S$.
@@
An intensity-positional allocation is Pareto-efficient.
@@
The sets of intensity-efficient, intensity-positional and Borda allocations are generally distinct.
That an intensity-positional allocation may not be intensity-efficient has already been shown in Example 5. The converse implication can also be seen by modifying the profile of Example 3 (Table 3) so that $s_{2}(a_{1},a_{3})=5$, $s_{2}(a_{2},a_{4})=4$ and $s_{3}(a_{1},a_{2})=4$, $s_{3}(a_{2},a_{4})=3$, $s_{3}(a_{2},a_{3})=2$, $s_{3}(a_{3},a_{4})=1$. In this case, the set of intensity-efficient allocations contains three elements, one of which—namely, $(a_{3},a_{2},a_{1},a_{4})$—is not intensity-positional. As far as the comparison between intensity-positional and Borda allocations is concerned, in light of Example 5 it suffices to show that there are profiles where some allocations are of the former but not of the latter kind. To this end, let $n=3$ and suppose $s_{1}(a,c)=3$, $s_{1}(b,c)=2$, $s_{1}(a,b)=1$; $s_{2}(a,b)=3$, $s_{2}(a,c)=2$, $s_{2}(c,b)=1$; and $s_{3}(b,a)=3$, $s_{3}(c,a)=2$, $s_{3}(b,c)=1$. Allocation $(b,a,c)$ is intensity-positional but not Borda. $\square$
@@
6 Concluding Remarks
In this paper we study the classic one-to-one assignment problem in an environment where, in addition to standard ordinal preferences, agents are also assumed to have ordinal preference intensities. These are internally consistent comparisons between preference improvements that do not require any quantification of the kind that is implied by cardinal utility functions. Two interpersonal comparability assumptions for such ordinal intensities are introduced and motivated. Building on this analytical environment, which provides a novel and potentially fertile middle ground between purely ordinal and fully cardinal environments for the study of allocation problems, we introduce and analyze the concepts of intensity-efficient and intensity-positional allocations. While distinct and with different points of strength and weakness, they share the feature that they refine the set of Pareto-efficient allocations in ways that incorporate this cognitively less demanding information about agents’ intensities into the allocation process with the goal—shared by notions defined in the cardinal-welfarist tradition—of achieving greater distributive justice.